OpenStudy (anonymous):

The function f(t)=5*4^t-1 approximates the number of frogs in a park in year t. What does the value 4 particularly represented in this situation A. There were 4 frogs in the park in Year 1 B. The numbers of frogs quadruples every year C. The park removed the frogs after year 4 D. The frog population increases by 4 every year

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@Hero

OpenStudy (sleepyjess):

I'm not sure, @Frostbite or @welshfella might be able to help though

OpenStudy (welshfella):

hmm. year 1 there are 5*4^(1-1) = 5 frogs 2 5*4^(2-1) = 20 frogs 3 5*4^(3-1) = 80 frogs

OpenStudy (welshfella):

So which one do you think ?

OpenStudy (welshfella):

5 , 20 , 80

OpenStudy (welshfella):

you get the nest value by multiplying by 4, right?

OpenStudy (anonymous):

Okay...I think that it is D

OpenStudy (anonymous):

Sorry about it taking so long...my internet is slow

OpenStudy (welshfella):

no increasing by 4 means you ADD 4 each year

OpenStudy (welshfella):

- we are MULTIPLYING by 4

OpenStudy (anonymous):

Oh wait...it is B

OpenStudy (anonymous):

Because you are adding 4 time the amount of frogs each year

OpenStudy (welshfella):

yes quadruple means ' 4 times'.

OpenStudy (anonymous):

Thank you!!

OpenStudy (welshfella):

yw